# EvalLM: Interactive Evaluation of Large Language Model Prompts on User-Defined Criteria

### v1

@@ 7.2.4. Trust
While a few participants expressed that they might “rely too much on the system” (P3), we observed that participants held a healthy level of skepticism about the evaluation assistant—the mean rating for trust was 4.91 (SD=1.51) out of 7. Participants mentioned how they would check the evaluation explanations and highlights to verify that the evaluations were “reasonable” (P11). Adding on this, P9 mentioned, “I didn’t consider the system to be an automatic evaluator since I was still checking its evaluation results”—illustrating a human-AI collaborative workflow where the system evaluates and user verifies. Some participants also mentioned how their trust was moderated by verifying evaluations as the explanations could “give a reason that [they] did not consider to be important” (P12) or were “not aligned” (P4) with their thoughts.
Despite being skeptical of the evaluation assistant, multiple participants mentioned that they might trust it more than their own manual evaluations. Participants expressed that they were “not completely confident” (P2) in their evaluations as they “couldn’t look at the whole text” (P6) and mostly evaluated “based on feeling” (P12). Additionally, participants felt that they be biased towards their own prompt as they were “the person creating the prompt and also evaluating it” (P11) and considered that the automatic evaluations could be more correct as the “AI might have more background knowledge than [they] do” (P10). Further, P5 and P8 mentioned how they attempted to “not compromise or bias” (P8) the evaluations by using different phrasing in their prompts and criteria. Although these findings portray the usefulness of the automatic evaluations, they also suggest that there is a potential danger in designers becoming over reliant on these evaluations.
@@ 7.2.5. Perceived Workload
Overall, participants felt significant lower mental burden (Assist=3.92 ± 1.78, Manual=5.58 ± 1.31, z=2.50, p=0.01) and significant lower effort (Assist=3.50 ± 1.68, Manual=5.25 ± 1.48, z=5.50, p=0.04) in the Assist condition (Fig. 10). The overall subjective mental workload was also lower in the Assist condition, although with marginal significance (Assist=3.45 ± 1.19, Manual=4.48 ± 0.99, z=13.00, pp=0.08). As described in the findings above, this can be attributed to how the Assist condition facilitated every step of the evaluation-refinement process: ideating on how to revise prompts, verifying the effect of revisions on performance, thinking of what to evaluate, and describing criteria.
@@ 8. Discussion
In this paper, we present EvalLM, an interactive system that supports designers in testing and comparing prompts by evaluating them on user-defined criteria with the aid of an LLM-based evaluator. We believe that EvalLM can narrow the gap between the development and deployment of LLM-based applications by helping designers iterate on their prompts when they cannot recruit external evaluators or testers to provide feedback. While EvalLM focuses on facilitating prompt evaluation, we believe that it can be expanded to facilitate prompt refinement, and enhance the evaluation and refinement of LLMs’ themselves. In this section, we discuss the further potential of EvalLM and opportunities for future work.
@@ 8.3. Beyond Prompts to Models
EvalLM is also capable of evaluating and comparing models. Instead of only uploading input samples, the system allows developers to upload accompanying output pairs. With the proliferation of high-performing but smaller-scale LLMs (e.g., LLaMA (Touvron et al., 2023a, b)) and the introduction of parameter-efficient fine-tuning (PEFT) methods (Mangrulkar et al., 2022; Hu et al., 2021; Zhang et al., 2023), developers and researchers have started to fine-tune their own LLMs to overcome the restrictions of prompting (Diao et al., 2023; Cui et al., 2023; Wang et al., 2023b). As the efficiency of LLM training increases and its cost decreases, we may see the proliferation of LLMs for a wider array of use cases and applications. In this environment, practitioners could use EvalLM as a first check on the performance of their trained models or, more promisingly, as a validation stage while the model is training to verify progress and identify weaknesses. Furthermore, as suggested in the previous subsection, practitioners can design an effective set of evaluation criteria through EvalLM and then employ this to align the models with these criteria through RLAIF. Thus, beyond prompt design, we believe that our work can also support practitioners in the development of more context- and task-specific models.
@@ 8.4. Evaluation Landscape for Natural Language Generation
Traditionally, research in NLG measured progress on how models perform on general-purpose tasks (e.g., summarization (Nallapati et al., 2016), topical conversations (Gopalakrishnan et al., 2023)), where performance was measured through more general criteria (e.g., “coherency”, “relevance” (Zhong et al., 2022; Fabbri et al., 2021)). As models become more capable of performing specific and long-tail tasks, on the other hand, developers and researchers may evaluate models on more task-specific criteria. While this diversification of criteria could lead to a more comprehensive understanding of LLM performance (Gehrmann et al., 2023), it can also become more challenging to synthesize the results from different evaluations and compare model performance. However, as shown by our formative interviews and user study, most of these task-specific criteria are frequently subordinate to more general criteria—meaning that results on specific criteria can present insights about performance on general criteria. Future work could collect and organize criteria into hierarchies that can represent model performance at both fine-grained and coarse-grained levels to enable practitioners to make more informed model choices.
@@ 9. Conclusion
This paper presents EvalLM, a novel interactive system that supports prompt designers in refining LLM prompts by evaluating generated outputs on user-defined criteria. In the interface, the user composes pairs of prompts and a set of evaluation criteria, and then employs LLMs to both generate outputs with the prompts and evaluate the outputs on the criteria. Through the explanations provided by the LLM on its evaluation, the user can iteratively refine their prompts, by identifying where they lack, and their criteria, by identifying where they are unclear. In a comparative user study (N=12), we compared EvalLM to a condition where participants manually evaluated outputs. We observed that EvalLM allowed participants to easily verify the effects of their prompt revisions, identify possible directions for improvement, and to define criteria that were more adapted to specific contexts.
@@ Acknowledgements.
This work was supported by KAIST-NAVER Hypercreative AI Center. We thank all of our participants for engaging positively in our various studies. We also thank all of the members of KIXLAB for their helpful discussions and constructive feedback.References
@@ Appendix A EvalLM: Experiment Screen
Figure 11 shows the Experiment screen in EvalLM, which allows users to test and evaluate their prompts on a larger nubmer of samples and test the reliability of the evaluations.
@@ Appendix B Prompts
Below, the full prompts used in EvalLM, where blue text represents content that is programmatically filled in.
@@ Appendix C Technical Evaluation: Details
Person: Introverted friend, who dislikes public speaking Action: Volunteer as a guest speaker at a local event | |
@@
Character: Sheldon from ”The Big Bang Theory” Question: ”What is your opinion on hand dryers?” |
@@ Appendix D User Study: Survey Questions
The questions for self-confidence in designing and improving prompts were combined into one measure for self-confidence in prompting, and the questions for self-confidence in planning and perfoming evaluations were combined into one measure for self-confidence in evaluating. The question for match goal, think-through, and collaborative were adapted from Wu et al. (Wu et al., 2022).

### v2

@@ 7.2.4. Trust
While a few participants expressed that they might “rely too much on the system” (P3), we observed that participants held a healthy level of skepticism about the evaluation assistant—the mean rating for trust was 4.91 (SD=1.51) out of 7. Participants mentioned how they would check the evaluation explanations and highlights to verify that the evaluations were “reasonable” (P11). Adding on this, P9 mentioned, “I didn’t consider the system to be an automatic evaluator since I was still checking its evaluation results”. Participants’ comments illustrated a human-AI collaborative workflow where the LLM evaluated multiple samples and they only verified a subset of these evaluations—allowing them to evaluate performance at a larger scale with less effort. Specifically, most participants focused on verifying evaluations where their prompt lost, tied, or had improved.
Although we observed that participants grew more skeptical when evaluations “gave a reason that [they] did not consider to be important” (P12) or were “not aligned” (P4) with their thoughts, multiple participants mentioned that they might still trust it more than their own evaluations. Participants were “not completely confident” (P2) in their own evaluations as they “couldn’t look through the whole text” (P6) and mostly evaluated “based on feeling” (P12), but the LLM “might have more background knowledge than [they] do” (P10). Also, participants felt that they could be biased as they were “the person creating the prompt and also evaluating it” (P11). In fact, P5 and P8 even tried to “not compromise or bias” (P8) the evaluations by purposefully using different phrasing in prompts and criteria. Although these findings portray the usefulness of LLM evaluations, they also hint at the potential danger of overreliance.
@@ 7.2.5. Perceived Workload
Participants felt significant lower mental burden (Assist=3.92 ± 1.78, Manual=5.58 ± 1.31, z=2.50, p=0.01) and significant lower effort (Assist=3.50 ± 1.68, Manual=5.25 ± 1.48, z=5.50, p=0.04) in the Assist condition (Fig. 10). The overall perceived workload was also lower in Assist, although with marginal significance (Assist=3.45 ± 1.19, Manual=4.48 ± 0.99, z=13.00, pp=0.08). As described in our findings, this can be attributed to how the Assist condition facilitated every step of the evaluation-refinement process: ideating on how to revise prompts, verifying how revisions affected performance, thinking of what to evaluate, and describing criteria.
@@ 8. Discussion
In this paper, we present EvalLM, an interactive system that supports designers in testing and comparing prompts by evaluating them on user-defined criteria with the aid of an LLM-based evaluator. We believe that EvalLM can narrow the gap between the development and deployment of LLM-based applications by helping designers iterate on their prompts when they cannot recruit external evaluators or testers to provide feedback. While EvalLM focuses on facilitating prompt evaluation, we believe that it can be expanded to facilitate prompt refinement, and enhance the evaluation and refinement of LLMs themselves. In this section, we discuss the further potential of EvalLM and opportunities for future work.
@@ 8.3. Beyond Prompts to Models
EvalLM is also capable of evaluating and comparing models. Instead of only uploading input samples, the system allows developers to upload accompanying output pairs. With the proliferation of high-performing but smaller-scale LLMs (e.g., LLaMA (Touvron et al., 2023a, b)) and the introduction of parameter-efficient fine-tuning (PEFT) methods (Mangrulkar et al., 2022; Hu et al., 2021; Zhang et al., 2023), developers and researchers have started to fine-tune their own LLMs to overcome the restrictions of prompting (Diao et al., 2023; Cui et al., 2023; Wang et al., 2023b). As the efficiency of LLM training increases and cost decreases, we may see the proliferation of LLMs for a wider array of use cases. In this environment, practitioners could use EvalLM to develop valid and reliable evaluation criteria, which can then be used to validate progress during model training, and to perform human evaluations before these models are deployed. Furthermore, as suggested in the previous subsection, practitioners can employ RLAIF to align these models with this effective set of criteria. Thus, beyond prompt design, we believe that our work can also support practitioners in the development of more context- and task-specific models.
@@ 8.4. Evaluation Landscape for Natural Language Generation
Traditionally, research in NLG measured progress on how models perform on general-purpose tasks (e.g., summarization (Nallapati et al., 2016), topical conversations (Gopalakrishnan et al., 2023)) by measuring performance on more general criteria (e.g., “coherency”, “relevance” (Zhong et al., 2022; Fabbri et al., 2021)). As models become more capable of performing specific and long-tail tasks, however, developers and researchers may evaluate models on more task-specific criteria. While this diversification of criteria could lead to a more comprehensive understanding of LLM performance (Gehrmann et al., 2023), it can also become more challenging to compile results from different evaluations and compare model performance. However, as shown by our formative interviews and user study, most of these task-specific criteria are frequently subordinate to more general criteria—meaning that results on specific criteria can present insights about performance on general criteria. Future work could collect and organize criteria into hierarchies that can represent model performance at both fine-grained and coarse-grained levels to enable practitioners to make more informed model choices.
@@ 9. Conclusion
This paper presents EvalLM, a novel interactive system that supports prompt designers in refining LLM prompts by evaluating generated outputs on user-defined criteria. In the interface, the user composes pairs of prompts and a set of evaluation criteria, and then employs LLMs to both generate outputs with the prompts and evaluate the outputs on the criteria. Through the explanations provided by the LLM on its evaluation, the user can iteratively refine their prompts, by identifying where they lack, and their criteria, by identifying where they are unclear. In a comparative user study (N=12), we compared EvalLM to a condition where participants manually evaluated outputs. We observed that EvalLM allowed participants to easily verify the effects of their prompt revisions, identify possible directions for improvement, and define criteria that were more adapted to specific contexts.
@@ Acknowledgements.
This work was supported by NAVER-KAIST Hypercreative AI Center. This work was also supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2021-0-01347,Video Interaction Technologies Using Object-Oriented Video Modeling). We thank all of our participants for engaging positively in our various studies. We also thank all of the members of KIXLAB for their helpful discussions and constructive feedback.References
@@ Appendix A EvalLM: Experiment Screen
Figure 11 shows the Experiment screen in EvalLM, which allows users to test and evaluate their prompts on a larger number of samples and test the reliability of the evaluations.
@@ Appendix B Prompts
Below, the full prompts used in EvalLM, where blue text represents content that is programmatically filled in. Regarding the context length limit of the LLMs, all of the prompts are around 300-500 tokens in length and criteria created by participants in our study were, on average, approximately 30 tokens in length. Considering GPT-4’s context window of 8,000 tokens and an approximate length of 200 tokens for an average paragraph, this means that a relatively large number of criteria and relatively lengthy outputs can be evaluated and reviewed using our prompts.
@@ Appendix C Technical Evaluation: Details
Person: Introverted friend, who dislikes public speaking
@@
Character: Sheldon from ”The Big Bang Theory”
@@ Appendix D User Study: Survey Questions
The questions for self-confidence in designing and improving prompts were combined into one measure for self-confidence in prompting, and the questions for self-confidence in planning and performing evaluations were combined into one measure for self-confidence in evaluating. The questions for match goal, think-through, and collaborative were adapted from Wu et al. (Wu et al., 2022).